The average height of American men is 69" with standard deviation of 3" and is assumed to be normally distributed. (PLEASE SHOW WORK)

A. How tall must a mattress be so that only 1% of American men find it too short?

B. What if a clothing manufacturer wishes to produce pants of a length to accommodate the middle 95% of the population of American men. What are the heights/lengths of the middle 95%?
Answer #1

A)for top 1% ; from normal value table ; crtiical z =2.33

therefore corresponding value =mean+z*std deviation =69+2.33*3=75.99 ~ 76.0''

B) for middle 95% values ; crtiical z =-/+ 1.96

therefore corresponding values =69-/+1.96*3=63.12 to 74.88